In spring and summer of 2010 a record number of 30 million sockeye salmon spawned up the Fraser river . In 2009 there were only just over 1 million sockeye spawning. Quite naturally the price of sockeye fell from about $ 5.00 per fish to about $1 per fish .
1. Assuming that 30% of the sockeye were caught each year , calculate the price elasticity of sockeye salmon , using the average price and quantity of your calculation.
2. Do you think the local fishermen were happy with such enormous run of fish ?
How to solve this problem?
1
Expert's answer
2019-10-08T09:43:57-0400

1. Catch 2010: 30 * 0.3 = 9 million

2. Catch 2009: 1 * 0.3 = 0.3 million

3. Elasticity: = ((9-0.3) / 0.3) / ((5-1) / 1) = 7.25

4. Thus, a given product is absolutely elastic in price, which indicates a close relationship between changes in demand and price.

Naturally, for fishermen, subject to domestic consumption, this situation is not favorable, since much more needs to be done to get profit at the 2009 level. In order to remedy this situation, a significant portion of the catch needs to be exported.